Animax sells 150 000 units of dog food to retail stores. The cost to the company is $25 per unit. The cost of processing each order is $260 and the carrying cost is $4.50 per unit. Required:

Using the EOQ model, calculate the quantity the company should order at a time.


1
Expert's answer
2021-07-13T11:07:57-0400

D=15000 S=$260 H=$4.5

EOQ= (2*D*S/H)^0.5

EOQ=(2*15000*260/4.5)^0.5=4163